OpenStudy (anonymous):

Find the volume of the solid of revolution obtained by rotating the curve x = e^t, y = cos t, t E [0, π], around the x axis.

OpenStudy (anonymous):

yikes calc 2 no fun

OpenStudy (anonymous):

I'm confused...... do you mean x=e^y and y=cos(x) and x E[0, pi]? because the problem as it is says y and x are both functions of t

OpenStudy (anonymous):

x = e^t y = cost \[t\in [0,π]\] Answer is supose to be 3π(e^π-1)/5 Cant find how i get it on my own it seems, but im still trying Thanks for all help

OpenStudy (anonymous):

Oh.....your working with parametric equations

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

do you have the equation for volume? I don't see it on my calc two book?

OpenStudy (anonymous):

Cant find a good example, but maybe this site http://tutorial.math.lamar.edu/Classes/CalcI/VolumeWithRings.aspx